Find the unit vector that is perpendicular to vectors A=2i + j + k and B= -1 + 2j + k

1
Expert's answer
2022-04-07T14:35:07-0400

"\hat{n} =\cfrac{\overrightarrow{A} \times\overrightarrow{B}} {|\overrightarrow{A} \times\overrightarrow{B}|}"


"\overrightarrow{A} \times\overrightarrow{B}=\begin{vmatrix}
 \hat{i} & \hat{j} & \hat{k}\\
 2 & 1 & 1\\
 - 1 & 2 & 1
\end{vmatrix}="

"=\hat{i}(1\cdot1-1\cdot2) -\hat{j}(2\cdot1-1\cdot(-1))+\\
+\hat{k}(2\cdot2-1\cdot(-1))=\\
=-\hat{i}-3\hat{j}+5\hat{k}"


"|\overrightarrow{A} \times\overrightarrow{B}|=\sqrt{(-1)^2+(-3)^2+5^2} =\sqrt{35}"


"\hat{n} =\cfrac{-\hat{i}} {\sqrt{35}} -\cfrac{3\hat{j}}{\sqrt{35}}+\cfrac{5\hat{k}}{\sqrt{35}}."